40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Apr 11 02:00:21 2013, Jenne, Update, Locking, TRY gone??? 
    Reply  Sun Apr 14 15:03:17 2013, Manasa, Update, Locking, Fixing - progress 
       Reply  Sun Apr 14 17:30:14 2013, Manasa, Update, Locking, Fixed 
          Reply  Sun Apr 14 23:20:42 2013, Den, Update, Locking, Fixed 
             Reply  Thu Apr 18 21:12:56 2013, Manasa, Update, Locking, Fixed 
Message ID: 8452     Entry time: Sun Apr 14 15:03:17 2013     In reply to: 8438     Reply to this: 8453
Author: Manasa 
Type: Update 
Category: Locking 
Subject: Fixing - progress 

Quote:

TRY signals are all gone!  Both the PD and the camera show no signal.  I went down there to turn off the lights, and look to see what was up, and I don't see any obvious things blocking the beam path on the table.  However, Steve has experimentally bungeed the lids down, so I didn't open the box to really look to see what the story is.

Absent TRY, I redid the IFO alignment.  Yarm locked, so I assumed it was close enough.  I redid Xarm alignment pretty significantly.  Transmission was ~0.5, which I got up to ~0.85 (which isn't too bad, since the PMC transmission is 0.74 instead of the usual 0.83).  I then aligned MICH, and PRM.  After fixing up the BS alignment, the POP beam wasn't hitting the POP PD in the center any more.  I centered the beam on the PD, although as Gabriele pointed out to me a week or two ago, we really need to put a lens in front of POP, since the beam is so big.  We're never getting the full beam when the cavity flashes, which is not so good.

Den is still working on locking, so I'll let him write the main locking report for the night.

We see that the PRC carrier lock seems to be more stable when we lock MICH with +1 for ITMY and -1 for ITMX, and PRCL with -1 for both ITMs.  This indicates that we need to revisit the systematic problem with using the PRM oplev to balance the coils, since that oplev has a relatively wide opening angle.  I am working on how to do this.

I'm fixing the TRY path.

I misaligned PRM and restored ETMY; but did not see the Y arm flashing. I am going ahead and moving the optics to get Y arm flashing again.

The slider values on the medm screen before touching any of them (for the record):

       tt1        tt2      itmy        etmy
p    -1.3886    0.8443     0.9320    -3.2583
y     0.3249    1.1407    -0.2849    -0.2751
ELOG V3.1.3-